# Tokens for Learning, Tokens for Unlearning: Mitigating Membership Inference Attacks in Large Language Models via Dual-Purpose Training

### v1

@@ Abstract
Large language models (LLMs) have become the backbone of modern natural language processing but pose privacy concerns about leaking sensitive training data. Membership inference attacks (MIAs), which aim to infer whether a sample is included in a model’s training dataset, can serve as a foundation for broader privacy threats. Existing defenses designed for traditional classification models do not account for the sequential nature of text data. As a result, they either require significant computational resources or fail to effectively mitigate privacy risks in LLMs. In this work, we propose a lightweight yet effective empirical privacy defense for protecting training data of language modeling by leveraging the token-specific characteristics. By analyzing token dynamics during training, we propose a token selection strategy that categorizes tokens into hard tokens for learning and memorized tokens for unlearning. Subsequently, our training-phase defense optimizes a novel dual-purpose token-level loss to achieve a Pareto-optimal balance between utility and privacy. Extensive experiments demonstrate that our approach not only provides strong protection against MIAs but also improves language modeling performance by around 10% across various LLM architectures and datasets compared to the baselines.
@@
Toan Tran, Ruixuan Liu, Li Xiong Emory University {vtran29;rliu51;lxiong}@emory.edu
@@ 1 Introduction
Large language models (LLMs) have become the foundation of modern natural language processing with a wide range of applications in various domains Chang et al. (2024). The rapidly increasing deployment of LLMs raises serious concerns about the data privacy Yao et al. (2024). LLMs have been shown to memorize the training data which can be later extracted by adversaries Carlini et al. (2023). Membership inference attacks (MIAs) Shokri et al. (2017); Li et al. (2024a) aim to infer whether a sample is included in a model’s training data, serving as the foundation of broader privacy threats Carlini et al. (2021b).
Due to the importance of understanding and mitigating MIAs, a significant amount of research has been conducted to design MIA defenses Hu et al. (2022b). However, most defenses focus on general machine learning models for classification tasks and do not account for the sequential nature of text data, while advanced MIAs for LLMs have leveraged such property. For example, the series of Min-K works Zhang et al. (2025); Shi et al. (2024) use the token-level loss on outlier tokens and significantly enhance MIAs for LLMs. Thus, conventional data sanitization or regularization techniques have limited defense effectiveness Kandpal et al. (2022); Liu et al. (2024b). And even the classic differentially private (DP) training algorithm Abadi et al. (2016) provides a strong defense, this approach comes at the inevitable cost of increased computation and reduced utility Li et al. (2022a); Bu et al. (2023b), which may not be desirable when the model trainer serves as the defender.
In this paper, we propose a defense mechanism for membership inference attacks on LLMs – DuoLearn. A recent study Lin et al. (2024) reveals that using a carefully selected subset of tokens during training can match or even surpass the performance of using all tokens in language modeling. In the meantime, MIAs mainly exploit loss-based signals associated with a sample Mattern et al. (2023); Carlini et al. (2021a). We observe that during training, some tokens carry stronger MIA signals and make the sample more vulnerable to MIAs. Thus, we leverage such token sequence nature of LLMs and propose a dynamic token selection strategy during training to proactively identify and categorize tokens into hard tokens (those with high losses) and memorized tokens (those with strong signals for MIA risks). Accordingly, we design a dual-objective loss function that performs learning via gradient descent on the hard tokens and unlearning via gradient ascent on the memorized tokens simultaneously in one backward pass, which makes the model learn useful information but not memorize specific training samples. Our contributions can be summarized as follows:
@@ 2.1 MIAs on LLMs
Membership inference attacks are a crucial privacy threat to machine learning models. There are a significant number of MIAs proposed for traditional classification models Hu et al. (2022b). Shokri et al. (2017) introduce membership inference attacks via analyzing the prediction probability difference between the training and testing samples. Yeom et al. (2018) connects MIAs to the overfitting phenomenon and proposes to use cross entropy loss as an MIA signal. However, due to the significant differences between LLMs and traditional classification models, some of these attacks are not applicable to LLMs, while others, though feasible, do not yield high attack performance. Therefore, there are non-trivial efforts to design suitable MIAs for LLMs. Carlini et al. (2021a) calibrate the sample loss with zlib entropy and reference models. Mattern et al. (2023) generate synthetic neighboring samples for each target sample then calculate the loss difference between them as the MIA signal. Shi et al. (2024) consider only top K lowest token losses for the MIA signal, while Zhang et al. (2025) perform z-score normalization for token losses, using the token vocabulary’s mean and standard deviation, then select top K z-scores. Fu et al. (2024) prompts the target LLM to generate a dataset which is used to train a reference attack model. Duan et al. (2024); Puerto et al. (2025) conduct systematic evaluations of MIAs on the pretrained LLMs. Liu et al. (2024b) design a privacy backdoor that can increase the membership inference risks.
@@ 2.2 LLM Memorization
The billion-parameter scale enhances LLM capabilities but also magnifies the privacy concerns. Carlini et al. (2021a, 2023) demonstrate that LLMs can memorize parts of their training data. There is potential leakages of LLMs generating the training data when prompted appropriately. These are known as exact memorization which can be utilized by the adversaries to extract the exact training data. Nasr et al. (2025) demonstrated that the LLM safety alignment fails to mitigate the privacy risks. It is feasible to undo the safety alignment via fine tuning and the adversaries can prompt the LLM to generate its training data.
@@ 2.3 Defenses Against MIAs
Overfitting is the root of membership inference risks Shokri et al. (2017). There are several works that proposed regularization techniques for traditional classification models such as weight decay and dropout Srivastava et al. (2014). While these regularization methods effectively reduces the membership inference risks in the traditional classification models Song and Mittal (2021), they are not sufficient to prevent memorization in LLMs Tirumala et al. (2022); Lee et al. (2022). Nasr et al. (2018) employ adversarial training. Tang et al. (2022) propose an ensemble architecture of models. These approaches are not practical for LLMs due to the expensive computing cost.
Generally, in the context of LLMs, there are still limited number of works on defense mechanisms against MIAs and memorization. There are two main approaches: sanitize training data and differential privacy (DP). Pilán et al. (2022) propose a practical method to protect Personally Identifiable Information (PII) by detecting and replacing PII with anonymized tokens. Shi et al. (2022) sanitize the PII tokens and pretrain on the sanitized data before conducting DP based fine-tuning on the original data. Lukas et al. (2023) demonstrates the effectiveness of sentence-level DP in mitigating the risks of leaking PII. These PII protection methods are effective but may not be sufficient to protect against MIAs because for each sample, the number of PII tokens is usually small Li et al. (2024b). Liu et al. (2024a) propose a method to perturb the training texts by leveraging memorization triggers that can effectively protect a small fraction of the training data against MIAs. Deduplicating the training corpus can reduce the risks of MIAs but not entirely eliminate them Kandpal et al. (2022).
The second popular approach conducts training/fine-tuning with Differentially-Private Stochastic Gradient Descent (DPSGD). Li et al. (2022b); Yu et al. (2022) show LLMs are strong differentially private learners. There are also a few works that aim to improve the DP training efficiency such as memory Bu et al. (2023b) and distributed training Bu et al. (2023a). DP training/fine-tuning usually offers strong privacy protection for LLMs Li et al. (2024b); Amit et al. (2024). Lowy et al. (2024) theoretically prove DP with a loose privacy budget can defend against MIAs. Despite efforts to improve the computing efficiency of DPSGD, differential privacy inherently introduces computational overhead, architectural constraints, and significant utility trade-off at scale Bu et al. (2024). To address the computational overhead and utility tradeoff of using DP on LLMs, Hans et al. (2024) proposes a non-DP practical masking mechanism, called Goldfish, that performs pseudo-random token masking for loss calculation to prevent memorization.
@@ 3 How Do Tokens Contribute to Membership Inference Risks?
Compared to conventional classification problems, membership inference attacks in language modeling have significant differences. In particular, each query in traditional classification models requires only one prediction. On the other hand, each query to language models involves multiple token predictions due to the sequential nature of text. This difference yields a question that how tokens contribute to overall sample-level membership inference risks. To answer this question, we perform a token-level analysis of membership inference risks. We calculate the MIA signal for each token as its prediction loss calibrated by a reference model Carlini et al. (2021a). A smaller signal value indicates that the model has a significantly higher confidence than other reference model on predicting the token.
@@
Figure 1 (left) illustrates the histogram of MIA signal values of tokens of a sample (see Figure 8 in Appendix 9 for additional histograms). The non-member sample distribution centers around zero, while the member sample skews to the negative side. Consequently, the average aggregated MIA signal is below zero for members but around zero for non-members, leading to membership inference risks. Moreover, the MIA signal values vary for different tokens, so some tokens contribute more to the membership inference risks than the others. Figure 1 (right) illustrates the MIA signal ranking of tokens of a member sample over training steps (see Figure 9 in Appendix 9 for additional samples). There is a complex changing dynamic in ranking between tokens before it becomes more stable at the end when the training converges. Overall, the analysis suggests that the sample-level membership inference risk in language modeling stem from the cumulative effect of many tokens. This poses challenges for defense methods that need token-level granularity to isolate and mitigate specific sources of leakage. Additionally, it is non-trivial to develop a defense method that widely affects a large number of tokens without disrupting the complex token dependencies essential for model utility.
@@ 4 Proposed Methodology – DuoLearn
Motivated by the analysis, we propose DuoLearn– a training framework that dynamically identifies hard tokens (those with higher calibrated losses) for learning and memorized tokens (those with strong MIA signals) for unlearning simultaneously. This way, the model learns useful information without memorizing specific training samples.
@@
Reference Modeling. Reference model ($\theta_{ref}$) shares an identical architecture with the training model ($\theta$). We fine-tune a reference model on a small portion of the original dataset (denoted as $\mathcal{T}_{\text{aux}}$) that can reflect the desired data distribution by standard causal language modeling (CLM), i.e., implementing next-token-prediction cross entropy loss ($\mathcal{L}_{CE}$):
@@
Token Selection. As shown in the previous analysis on LLM generalization by Lin et al. (2024) and ours on membership inference risks, tokens contribute differently. Considering all tokens equally as the standard causal language modeling is not optimal since it can lead to memorization on some tokens and amplify the memorization over training epochs. DuoLearn defines two sets of tokens: hard tokens ($\mathcal{T}_{h}$) and memorized tokens ($\mathcal{T}_{m}$). Hard tokens are the tokens that the current training models ($\theta$) have difficulty predicting, while memorized tokens are the tokens that the model has already memorized. To identify these two sets of tokens, we propose a token selection mechanism based on the prediction loss of each token calibrated by the reference model. We implement the score $s(t_{i})$ for each token $t_{i}$ which is the difference between the cross-entropy loss of the training model and the reference model, as used in previous works Lin et al. (2024); Mindermann et al. (2022):
$s(t_{i})=\log P(t_{i}|t_{<i};\theta_{ref})-\log P(t_{i}|t_{<i};\theta).$
@@
Dual-Purpose Loss. We introduce a dual-purpose loss function designed to improve model performance on hard tokens while mitigating overfitting on memorized tokens. This loss function combines two components: the learning loss and the unlearning loss. The learning loss is the standard causal language modeling (CLM) loss applied to the hard tokens $\mathcal{T}_{h}$. The unlearning loss, in contrast, is the negative CLM loss applied to the memorized tokens $\mathcal{T}_{m}$, effectively performing gradient ascent. The dual-purpose loss is defined as follows, where $\alpha>0$ is a hyper-parameter that balances the learning and unlearning losses:
@@ 5.1 Experiment Settings
$f_{\text{Loss}}(x)=\mathcal{L}_{\text{CE}}(\theta;x)$
@@
$f_{\text{Ref}}(x)=\mathcal{L}_{\text{CE}}(\theta;x)-\mathcal{L}_{\text{CE}}(\theta_{\text{attack}};x)$
$\bullet$ Min-K Shi et al. (2024) leverages top K tokens that have the lowest loss values.
$f_{\text{min-K}}(x)=\frac{1}{|\text{min-K(x)}|}\sum_{t_{i}\in\text{min-K(x)}}-\log(P(t_{i}|t_{<i};\theta)$
@@
$f_{\text{zlib}}(x)=\mathcal{L}_{\text{CE}}(\theta;x)/\text{zlib}(x)$
Baselines. We present the results of four baselines. Base refers to the pretrained LLM without fine tuning. FT represents the standard causal language modeling without protection. Goldfish Hans et al. (2024) implements a masking mechanism. DPSGD Abadi et al. (2016); Yu et al. (2022) applies gradient clipping and injects noise to achieve sample-level differential privacy.
Implementation. We conduct full fine-tuning for GPT-2 and Pythia. For computing efficiency, Llama-2 fine-tuning is implemented using Low-Rank Adaptation (LoRA) Hu et al. (2022a) which leads to ~4.2M trainable parameters. Additionally, we use subsets of 3K samples to fine-tune the LLMs. We present other implementation details in Appendix C.1.

### v2

@@ Abstract
Large language models (LLMs) have become the backbone of modern natural language processing but pose privacy concerns about leaking sensitive training data. Membership inference attacks (MIAs), which aim to infer whether a sample is included in a model’s training dataset, can serve as a foundation for broader privacy threats. Existing defenses designed for traditional classification models do not account for the sequential nature of text data. As a result, they either require significant computational resources or fail to effectively mitigate privacy risks in LLMs. In this work, we propose DuoLearn, a lightweight yet effective empirical privacy defense for protecting training data of language models by leveraging token-specific characteristics. By analyzing token dynamics during training, we propose a token selection strategy that categorizes tokens into hard tokens for learning and memorized tokens for unlearning. Subsequently, our training-phase defense optimizes a novel dual-purpose token-level loss to achieve a Pareto-optimal balance between utility and privacy. Extensive experiments demonstrate that our approach not only provides strong protection against MIAs but also improves language modeling performance by around 10% across various LLM architectures and datasets compared to the baselines. [^1] The implementation code for DuoLearn is available at https://github.com/Emory-AIMS/duolearn
@@
Toan Tran, Ruixuan Liuthanks: Corresponding author: ruixuan.liu2@emory.edu, Li Xiong Emory University Atlanta, GA, USA {vtran29;ruixuan.liu2;lxiong}@emory.edu Appeared at ACL’25 (Findings)
@@ 1 Introduction
Large language models (LLMs) have become the foundation of modern natural language processing with a wide range of applications in various domains Chang et al. (2024). The rapidly increasing deployment of LLMs raises serious concerns about data privacy Yao et al. (2024). LLMs have been shown to memorize the training data which can be later extracted by adversaries Carlini et al. (2023). Membership inference attacks (MIAs) Shokri et al. (2017); Li et al. (2024a) aim to infer whether a sample is included in a model’s training data, serving as the foundation of broader privacy threats Carlini et al. (2021b).
Due to the importance of understanding and mitigating MIAs, a significant amount of research has been conducted to design MIA defenses Hu et al. (2022b). However, most defenses focus on general machine learning models for classification tasks and do not account for the sequential nature of text data, while advanced MIAs for LLMs have leveraged this property. For example, the series of Min-K works Zhang et al. (2025); Shi et al. (2024) uses the token-level loss on outlier tokens and significantly enhance MIAs for LLMs. Thus, conventional data sanitization or regularization techniques have limited defense effectiveness Kandpal et al. (2022); Liu et al. (2024b). Even though the classic differentially private (DP) training algorithm Abadi et al. (2016) provides a strong defense, this approach comes at the inevitable cost of increased computation and reduced utility Li et al. (2022a); Bu et al. (2023b), which may not be desirable when the model trainer serves as the defender.
In this paper, we propose a defense mechanism for membership inference attacks on LLMs – DuoLearn. A recent study Lin et al. (2024) reveals that using a carefully selected subset of tokens during training can match or even surpass the performance of using all tokens in language modeling. In the meantime, MIAs mainly exploit loss-based signals associated with a sample Mattern et al. (2023); Carlini et al. (2021a). We observe that during training, some certain tokens carry stronger MIA signals than others, making the sample vulnerable to MIAs. Thus, we leverage the token sequence nature of LLMs and propose a dynamic token selection strategy during training to proactively identify and categorize tokens into hard tokens (those the model struggles with) and memorized tokens (those with strong signals for MIA risks). Accordingly, we design a dual-objective loss function that performs learning via gradient descent on the hard tokens and unlearning via gradient ascent on the memorized tokens simultaneously in one backward pass, which makes the model learn useful information but not memorize specific training samples. Our contributions can be summarized as follows:
@@ 2.1 MIAs on LLMs
Membership inference attacks are a crucial privacy threat to machine learning models. There are a significant number of MIAs proposed for traditional classification models Hu et al. (2022b). Shokri et al. (2017) introduce MIAs via training a binary classification model over behaviors collected from shadow models. Yeom et al. (2018) connect MIAs to the overfitting phenomenon and propose to use cross entropy loss as an MIA signal. However, due to the significant differences between LLMs and traditional classification models, some of these attacks are not applicable to LLMs, while others, though feasible, have limited attack performance. Therefore, there are non-trivial efforts to design suitable MIAs for LLMs. Carlini et al. (2021a) calibrate the sample loss with zlib entropy and reference models. Mattern et al. (2023) generate synthetic neighboring samples for each target sample then calibrate the loss of the target sample with the averaged loss of its neighboring samples as the MIA signal. Shi et al. (2024) consider only top $K$ lowest token losses for the MIA signal, while Zhang et al. (2025) perform z-score normalization for token losses, using the token vocabulary’s mean and standard deviation, then select top $K$ z-scores. Fu et al. (2024) prompt the target LLM to generate a dataset which is used to train a reference attack model. Duan et al. (2024) conduct systematic evaluations of MIAs on the pretrained LLMs. Hayes et al. (2025) scale reference-based MIAs on large-scale pretraining settings. Puerto et al. (2025) consider various scales of membership from sentences to collections of documents. Liu et al. (2024b) design a privacy backdoor that can increase the membership inference risks. Feng et al. (2025) investigate MIAs on preference data used for post-training alignment.
@@ 2.2 LLM Memorization
The billion-parameter scale enhances LLM capabilities but also magnifies the privacy concerns. Carlini et al. (2021a, 2023) demonstrate that LLMs can memorize parts of their training data. There is a risk that LLMs may generate the training data when prompted appropriately. These are known as exact memorization which can be utilized by the adversaries to extract the exact training data. Nasr et al. (2025) demonstrated that the LLM safety alignment fails to mitigate the privacy risks. It is feasible to undo the safety alignment via fine-tuning and the adversaries can prompt the LLM to generate its training data.
@@ 2.3 Defenses Against MIAs
Overfitting is the root of membership inference risks Shokri et al. (2017). While regularization methods such as weight decay and dropout Srivastava et al. (2014) mitigates overfitting and slightly reduces the membership inference risks in the traditional classification models Song and Mittal (2021), they are not sufficient to prevent memorization in LLMs Tirumala et al. (2022); Lee et al. (2022). Other defenses which leverage adversarial training Nasr et al. (2018) or ensemble architecture of models Tang et al. (2022) are infeasible for LLMs due to the expensive computing cost.
Generally, in the context of LLMs, there are still limited number of works on defense mechanisms against MIAs and memorization. There are two main approaches: sanitizing training data and differential privacy (DP). Pilán et al. (2022) propose a practical method to protect Personally Identifiable Information (PII) by detecting and replacing PII with anonymized tokens. Shi et al. (2022) sanitize the PII tokens and pretrain on the sanitized data before conducting DP based fine-tuning on the original data. Lukas et al. (2023) demonstrate the effectiveness of sentence-level DP in mitigating the risks of leaking PII. These PII protection methods are effective but may not be sufficient to protect against MIAs because for each sample, the number of PII tokens is usually small Li et al. (2024b). Liu et al. (2024a) propose a method to perturb the training texts by leveraging memorization triggers that can effectively protect a small fraction of the training data against MIAs. Deduplicating the training corpus can reduce the risks of MIAs but not entirely eliminate them Kandpal et al. (2022).
The second popular approach conducts training/fine-tuning with Differentially-Private Stochastic Gradient Descent (DPSGD). Li et al. (2022b); Yu et al. (2022); Li et al. (2024b); Amit et al. (2024) show LLMs are strong differentially private learners. Lowy et al. (2024) theoretically prove that DP with a loose privacy budget can defend against MIAs. Despite efforts to improve the computing efficiency of DPSGD Bu et al. (2023b), DP inherently introduces computational overhead, architectural constraints, and significant utility trade-off at scale. McKenna et al. (2025) explore the scaling laws of DP LLMs and reveal challenges especially about the optimal training data size. To avoid the computational overhead and utility tradeoff of using DP on LLMs, Hans et al. (2024) proposes a non-DP practical masking mechanism, called Goldfish, that performs pseudo-random token masking for loss calculation to prevent memorization. Our method is also an empirical defense in the training stage.
@@ 3 How Do Tokens Contribute to Membership Inference Risks?
Compared to conventional classification problems, membership inference attacks in language modeling have significant differences. In particular, each query in traditional classification models requires only one prediction. On the other hand, each query to language models involves multiple token predictions due to the sequential nature of text. This difference yields a question that how tokens contribute to overall sample-level membership inference risks. To answer this question, we perform a token-level analysis of membership inference risks. We calculate the MIA signal for each token as its prediction loss calibrated by a reference model Carlini et al. (2021a). A smaller signal value indicates that the model has a significantly higher confidence than the reference model on predicting the token.
@@
Figure 1 (left) illustrates the histogram of MIA signal values for the tokens of a member sample and a non-member sample (see Figure 9 in Appendix 10 for additional histograms). The non-member sample distribution centers around zero, while the member sample skews to the negative side. Consequently, the average aggregated MIA signal is below zero for the member but around zero for the non-member, leading to membership inference risks. Moreover, the MIA signal values vary for different tokens, so some tokens contribute more to the membership inference risks than the others. Figure 1 (right) illustrates the MIA signal ranking of tokens of a member sample over training steps (see Figure 10 in Appendix 10 for additional samples). There is a complex changing dynamic in ranking between tokens before it becomes more stable at the end when the training converges. Overall, the analysis suggests that the sample-level membership inference risk in language modeling stem from the cumulative effect of many tokens. This poses challenges for defense methods, as they need token-level granularity to isolate and mitigate specific sources of leakage. Additionally, it is non-trivial to develop a defense method that widely affects a large number of tokens without disrupting the complex token dependencies essential for model utility.
@@ 4 Proposed Methodology – DuoLearn
Motivated by the analysis, we propose DuoLearn– a training framework that dynamically identifies hard tokens (those with higher calibrated losses) for learning and memorized tokens (those with lower calibrated loss or stronger MIA signals) for unlearning simultaneously. This way, the model learns useful information without memorizing specific training samples.
@@
Reference Modeling. Reference model ($\theta_{\text{ref}}$) shares an identical architecture with the training model ($\theta$). We fine-tune a reference model on a small portion of the original dataset (denoted as $\mathcal{T}_{\text{aux}}$) that can reflect the desired data distribution by standard causal language modeling (CLM), i.e., implementing next-token-prediction cross entropy loss ($\mathcal{L}_{CE}$):
@@
Token Selection. As our previous analysis, tokens contribute differently in membership inference risks. Thus, considering all tokens equally is not optimal for privacy defense with respect to the utility and privacy trade-off. DuoLearn defines two sets of tokens: hard tokens ($\mathcal{T}_{h}$) and memorized tokens ($\mathcal{T}_{m}$). Hard tokens are the tokens that the current training models ($\theta$) have difficulty predicting, while memorized tokens are the tokens that the model has already memorized. To identify these two sets of tokens, we propose a token selection mechanism based on the prediction loss of each token calibrated by the reference model. We implement the score $s(t_{i})$ for each token $t_{i}$ which is the difference between the cross-entropy loss of the training model and the reference model, as used in previous works Lin et al. (2024); Mindermann et al. (2022):
$s(t_{i})=\log P(t_{i}|t_{<i};\theta_{\text{ref}})-\log P(t_{i}|t_{<i};\theta).$
@@
Dual-Purpose Loss. We introduce a dual-purpose loss function designed to improve model performance on hard tokens while mitigating overfitting on memorized tokens. This loss function combines two components: the learning loss and the unlearning loss. The learning loss is the standard causal language modeling (CLM) loss applied to the hard tokens $\mathcal{T}_{h}$. The unlearning loss, in contrast, is the negative CLM loss applied to the memorized tokens $\mathcal{T}_{m}$, effectively performing gradient ascent. The dual-purpose loss is defined as follows, where $\alpha>0$ is a hyper-parameter that balances the learning and unlearning losses
@@ 5.1 Experiment Settings
$f_{\text{Loss}}(x)=-\mathcal{L}_{\text{CE}}(\theta;x)$
@@
$f_{\text{Ref}}(x)=\mathcal{L}_{\text{CE}}(\theta_{\text{ref}};x)-\mathcal{L}_{\text{CE}}(\theta;x)$
$\bullet$ Min-K Shi et al. (2024) leverages top K tokens that have the lowest probabilities.
$f_{\text{min-K}}(x)=\frac{1}{|\text{min-K(x)}|}\sum_{t_{i}\in\text{min-K(x)}}\log P(t_{i}|t_{<i};\theta)$
@@
$f_{\text{zlib}}(x)=-\mathcal{L}_{\text{CE}}(\theta;x)/\text{zlib}(x)$
Baselines. We present the results of four baselines. Base refers to the pretrained LLM without fine-tuning. FT represents the standard causal language modeling without protection. Goldfish Hans et al. (2024) implements a masking mechanism. DPSGD Abadi et al. (2016); Yu et al. (2022) applies gradient clipping and injects noise to achieve sample-level differential privacy.
Implementation. We conduct full fine-tuning for GPT-2 and Pythia. For computing efficiency, Llama-2 fine-tuning is implemented using Low-Rank Adaptation (LoRA) Hu et al. (2022a) which leads to ~4.2M trainable parameters. Additionally, we use subsets of 3K samples to fine-tune the LLMs. The data used to train DuoLearn’s reference model is disjoint from either the target model’s or the reference attack model’s training data. We present additional implementation details in Appendix C.1.